OpenStudy (anonymous):

Is my answer right? Yes or no? Thanks! Solve. 6x2 = 12x - 24 2 _+_ 3i sq root 2 over 2

OpenStudy (anonymous):



OpenStudy (cwrw238):

6(x^2 - 2x + 4) = 0 x = 2 +/- sqrt(4 - 4*1*4) / 2 = 1 +/- sqrt12i / 2 = 1 +/- (sqrt 4*3)i / 2 = 1 +/- i sqrt3